OpenStudy (kamille):

13. A circular bacterial chromosome is 1.28 mm long. Under optimal conditions the chromosome is replicated in 40 minutes. a) Under optimal conditions, assuming DNA replication is the same as in E. coli, what is the distance traversed by one replication fork in 1 minute? b) Given that the length occupied by 1 bp is 0.34 nm how many nucleotides are incorporated in 1 minute in one replication fork? Explain your workings.

OpenStudy (aaronq):

b) use a ratio \(\dfrac{1.28~mm}{40 ~min}=\dfrac{x~mm}{60~min} \rightarrow x=1.92~mm\) c) 1 mm= \(1.0*10^6~nm\), thus 1.92 mm = \(1.92*10^6~nm\) number of nucleotides =\(\dfrac{ 1.92*10^6~nm}{0.34~nm/nucleotide}=5647058.82 =5.7*10^6~nucleotides\)

OpenStudy (anonymous):

Which piece of evidence did Alfred Wegener use to develop the theory of continental drift?

OpenStudy (aaronq):

?

OpenStudy (anonymous):

A. undersea mountain ranges B. similar fossils on different continents C. basalts on the ocean floor D. paleomagnetism on the seafloor

OpenStudy (aaronq):

open your own question, don't post yours on other people.